OpenStudy (anonymous):

Which expression is equal to sec x° multiplied by cot x°? hypotenuse ÷ adjacent adjacent ÷ hypotenuse opposite ÷ adjacent hypotenuse ÷ opposite I think its the last one

OpenStudy (kirbykirby):

\[ \sec x = \frac{1}{\cos x}\] \[ \cot x = \frac{\cos x}{\sin x}\] Can you see how this simplifies :) Then just use the definition of your basic trig identities

OpenStudy (anonymous):

so its not the last one? haha sorry im so confused

OpenStudy (kirbykirby):

Oh sorry I didn't read the last sentence.. I thought all the stuff below the question were choices :) Indeed, you are right!

OpenStudy (anonymous):

haha ok thank you! I was just double checking it! thank you very much

OpenStudy (anonymous):

do you think you can help me with one more?

OpenStudy (kirbykirby):

sure

OpenStudy (anonymous):

OK thank you! this is the question! Find the measure of angle y. Round your answer to the nearest hundredth. (please type the numerical answer only)

OpenStudy (kirbykirby):

You are given the opposite length and the hypotenuse... so you can see that: \[\sin y=\frac{4}{20} \]

OpenStudy (kirbykirby):

Now you can solve for \(y\) by using the inverse operation \(\arcsin\), or also known as \(\sin^{-1}\)

OpenStudy (anonymous):

I have no clue on how to do that. im sorry could you plz walk me through it. I didnt understand anything in this lesson and my teacher is super busy

OpenStudy (kirbykirby):

Ok. So the way to cancel out the " \(\sin\)" is by using operation called \(\arcsin \) , but your teacher may have used the notation \(\sin^{-1}\) instead. It's the same thing though :) So: \[ \sin y = \frac{4}{20}\]\[ y=\arcsin\left(\frac{4}{20} \right)\] That last line is the same as writing \[y=\sin^{-1}\left(\frac{4}{20} \right) \]... now to find this quantity, you need to punch it into a calculator

OpenStudy (anonymous):

ummmm i'm getting decimals and I have the feeling thats totally wrong. how do I put it in the calculator I wanna make sure i'm doing it right

OpenStudy (kirbykirby):

Not sure which one you are using, but you should have a "shift" or "2nd function" button. Press that and then press the "\(\sin\)" button

OpenStudy (anonymous):

ok. I keep getting 0.2

OpenStudy (anonymous):

oh wait ok now I got 5 haha I put it in wrong.

OpenStudy (kirbykirby):

That's good if you need your answer in radians. If you need it in degrees, you should change the setting on your calculator

OpenStudy (kirbykirby):

Hm it shouldn't be 5 degrees. Be careful when converting from radian to degrees (or just set your calculator in degree mode and do the calculation again)

OpenStudy (anonymous):

OK now i'm confused. I'm using the only calculator I have which is my phone and it doesn't have that. ugh this is frustrating

OpenStudy (anonymous):

I'm sorry I'm probably annoying the crap outta you. I'm terrible at math

OpenStudy (kirbykirby):

Oh well o.o Um I suggest getting a scientific calculator... you would probably need one for your exams I believe

OpenStudy (kirbykirby):

No annoyance lol, don't worry.

OpenStudy (kirbykirby):

For now you can try this: http://www.wolframalpha.com/input/?i=arcsin%284%2F20%29 It gives you the numerical value in radians and in degrees